# On the Euclidean Distance Degree of Quadratic Two-Neuron Neural Networks

### v2

@@ 1 Introduction
The algebraic geometry of neurovarieties has attracted increasing attention in recent years. This is due to the central role played by algebraic and semi-algebraic models in Deep Learning, supported by density and universal approximation results [SM17, SM18], by the existence of global invariants describing the expressive power of architectures [KTB19], by the relative tractability of the inevitable singularities these varieties exhibit — that are more delicate to work with in a differential-geometric setting [Wat09] — and by the appearance of programmatic works such as [MSMTK25]. A central invariant in Machine Learning is the $\mathrm{ED}$-degree, which measures the number of functions within a model class that optimally fit the training data. This invariant depends sensitively on the choice of a metric on the ambient space; however, for a sufficiently general scalar product its value stabilizes. This general value is shown to be a projective invariant and it is given by the sum of the polar classes of the variety, which we call the virtual $\mathrm{ED}$-degree of the variety (see [DHOST] for a general treatment of this topic). In this paper we study the virtual $\mathrm{ED}$-invariant for the neurovarieties associated to a shallow model with two neurons and scalar output and we use Intersection Theory to show that this value is a stably polynomial function on the dimension of the input space.
@@ 3 Neurovarieties with scalar output and the $\mathrm{ED}$-degree
We introduce here the central notions of the paper. The definitions and results in this section are stated in a slightly greater generality than what will be strictly needed later on. This choice is motivated by the fact that we think such a level of generality provides a clearer geometric picture and helps place our main result in the broader context of Euclidean distance degree theory. Further details can be found in [KTB19, KMMT22, Sha24].
@@ 3.4 The virtual $\mathrm{ED}$-degree
This is the point of view adopted in [DHOST]. In particular, if we denote with $\mathbb{P}\left(\mathcal{E}^{(u)}_{X}\right)$ the projectivization of the $\mathrm{ED}$-correspondence for a generic $u\in\mathbb{C}^{N+1}$ we have
@@ 4.1 Polar degrees, Chern–Mather classes and Nash blow-up of $\mathcal{M}$
When descending to the projective variety two first-order deformations $A+\varepsilon B$ and $A+\varepsilon B^{\prime}$ define the same tangent direction at $\left[A\right]\in\mathcal{M}$ if and only if they differ by the action of $\mathbb{C}_{\varepsilon}^{\ast}$, that is
$A+\varepsilon B^{\prime}=\left(1+\varepsilon\lambda\right)\left(A+\varepsilon B\right)\qquad\text{in }\mathrm{Sym}^{2}V^{\ast}\otimes\mathbb{C}_{\varepsilon},$
for some $\lambda\in\mathbb{C}$ or, equivalently, $B^{\prime}-B=\lambda A$. Therefore the embedded projective tangent space at $\left[A\right]$ is
$T_{\left[A\right]}\mathcal{M}\simeq\frac{T_{A}C\left(\mathcal{M}\right)}{\left\langle A\right\rangle}.$ (4.6)

### v3

@@ 1 Introduction
The algebraic geometry of neurovarieties has attracted increasing attention in recent years. This is due to the central role played by algebraic and semi-algebraic models in Deep Learning, supported by density and universal approximation results [SM17, SM18], by the existence of global invariants describing the expressive power of architectures [KTB19], by the relative tractability of the inevitable singularities these varieties exhibit — that are more delicate to work with in a differential-geometric setting [Wat09] — and by the appearance of programmatic works such as [MSMTK25]. A central invariant in Machine Learning is the $\mathrm{ED}$-degree, which measures the number of functions within a model class that optimally fit the training data. This invariant depends sensitively on the choice of a metric on the ambient space; however, for a sufficiently general scalar product its value stabilizes. This general value is shown to be a projective invariant and it is given by the sum of the polar classes of the variety, which we call the virtual $\mathrm{ED}$-degree of the variety $v\mathrm{ED}$ (see [DHOST] for a general treatment of this topic). In this paper we study the virtual $\mathrm{ED}$-invariant, that for the neurovarieties associated to a shallow model with two neurons and scalar output and we use Intersection Theory to show that this value is a stably polynomial function on the dimension of the input space. Closely related models were studied in [KLW24], where the authors allow the output dimension to vary and, by exploiting explicit formulas relating the virtual $\mathrm{ED}$-degree to intersection matrices on suitable Grassmannians, obtain closed expressions for $v\mathrm{ED}$.
@@ 3 Neurovarieties with scalar output and the $\mathrm{ED}$-degree
We introduce here the central notions of the paper. The definitions and results in this section are stated in a slightly greater generality than what will be strictly needed later on. This choice is motivated by the fact that we think such a level of generality provides a clearer geometric picture and helps place our main result in the broader context of Euclidean distance degree theory. Further details can be found in [KTB19, KMMT22, Sha24]. See also [KLW24] for a comprehensive introduction to polynomial neural networks and numerous examples.
@@ 3.4 The virtual $\mathrm{ED}$-degree
This is the point of view adopted in [DHOST]. In particular, if we denote with $\mathbb{P}\left(\mathcal{E}^{(u)}_{X,Q}\right)$ the projectivization of the $\mathrm{ED}$-correspondence for a generic $u\in\mathbb{C}^{N+1}$ we have
@@ 4.1 Polar degrees, Chern–Mather classes and Nash blow-up of $\mathcal{M}$
When passing to the projective variety, tangent directions at a point $\left[A\right]\in\mathcal{M}\subset\mathbb{P}\left(\mathrm{Sym}^{2}V^{\ast}\right)$ are described in terms of first-order deformations over the dual numbers $\mathbb{C}_{\varepsilon}$. By definition, a $\mathbb{C}_{\varepsilon}$–point of $\mathbb{P}\left(\mathrm{Sym}^{2}V^{\ast}\right)$ corresponds to a rank-one direct summand of $\mathrm{Sym}^{2}V^{\ast}\otimes\mathbb{C}_{\varepsilon}$ whose reduction modulo $\left(\varepsilon\right)$ is $\langle A\rangle$. Accordingly, two first-order deformations define the same tangent direction at $\left[A\right]$ if and only if they determine the same rank-one $\mathbb{C}_{\varepsilon}$-submodule of $\mathrm{Sym}^{2}V^{\ast}\otimes\mathbb{C}_{\varepsilon}$. After fixing a representative whose reduction modulo $\left(\varepsilon\right)$ is $A$, this equivalence identifies deformations whose infinitesimal parts differ by a multiple of $A$. In other words, the tangent direction only depends on the class of $B$ in $\mathrm{Sym}^{2}V^{\ast}/\langle A\rangle$. Therefore, the embedded projective tangent space at $\left[A\right]$ is canonically identified with
$T_{\left[A\right]}\mathcal{M}\simeq\frac{T_{A}C\left(\mathcal{M}\right)}{\langle A\rangle}.$ (4.6)